the average of two numbers is 11. Their difference is 3 more than that. Find the numbers
Answer by ewatrrr(24785)


Hi
Let x and y represent the numbers
Question states***
(x+y)/2 = 11 OR x+y = 22
x-y = 14
x+y = 22
x-y = 14
2x = 36
x = 18 and y = 4 (y = 18-14)
CHECKING our Answer***
(18+4)/2 = 11